OpenStudy (anonymous):

the steps of how to solve 8 = sqrt(2^(4x + 8))

OpenStudy (anonymous):

8 = sqrt(2^(4x + 8)) 8^2 = sqrt(2^(4x + 8))^2 64 = 2^(4x + 8) 2^6 = 2^(4x + 8) 6 = 4x + 8 6 - 8 = 4x + 8 - 8 -2 = 4x 4x / 4 = -2 / 4 x = -1 / 2

OpenStudy (anonymous):

how would you explain that in words

OpenStudy (anonymous):

the easiest way is in numbers

OpenStudy (anonymous):

my assignment asked to explain in sentences thats why i asked

OpenStudy (anonymous):

oh ok